OpenStudy (zeesbrat3):

If f(1) = 5 and f prime of x equals the quotient of 6 and the quantity x squared plus 2, which of the following is the best approximation for f(1.03) using local linearization?

OpenStudy (zeesbrat3):

\[f(1) = 5\] \[f'(x) = \frac{ 6 }{ x^2 + 2}\]

OpenStudy (zeesbrat3):

@ganeshie8

OpenStudy (solomonzelman):

you first need to integrate the 6/(x²+2)

OpenStudy (solomonzelman):

you would need a trig substitution....

OpenStudy (zeesbrat3):

I haven't gotten to integration yet..

OpenStudy (solomonzelman):

oh, then just recongize the derivative I guess

ganeshie8 (ganeshie8):

\[f(1+0.03) = f(1) + f'(1)(1.03-1)\]

OpenStudy (solomonzelman):

I feel stupid now:)

OpenStudy (solomonzelman):

you don't even need to find f(x).

OpenStudy (zeesbrat3):

There is no defined f'(1) I thought?

ganeshie8 (ganeshie8):

Since it is an approximation, it makes more sense to use \(\approx\) \[f(1+0.03) \approx f(1) + f'(1)(1.03-1)\] simplify

OpenStudy (solomonzelman):

for f`(1) plug in 1 for x into the f'(x)

OpenStudy (solomonzelman):

(and it is always defined, because x²+6 is never=0, it is at least 6)

OpenStudy (zeesbrat3):

So \[5 + 2(.03)\]

ganeshie8 (ganeshie8):

Looks good! it simplifies further tho

OpenStudy (zeesbrat3):

To 5.06

ganeshie8 (ganeshie8):

Yep

OpenStudy (zeesbrat3):

Thank you!

OpenStudy (solomonzelman):

I wish one day I would be that good at math too.... :D f(1+0.03) was quite awesome, I will remeber this trick.

ganeshie8 (ganeshie8):

yw

OpenStudy (solomonzelman):

tnx ganesh.

ganeshie8 (ganeshie8):

np